A simple random sample of 15 people from a certain population has mean age of 25 with a standard deviation of 10. Can we conclude that the mean age of the population is younger than 25? Let alpha = 0.05

  • Formulate the hypothesis: the null hypothesis and the alternative hypothesis.
  • Set the significance level for a. type of test critical value of z/t
  • Compute the z-statistic or t-statistic
  • Decision: compare the computed value of z or t with the Critical value.
  • Conclusion

can anyone help me with this? please :(


1
Expert's answer
2022-05-08T13:40:39-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge25"

"H_a:\mu<25"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


2. Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=14" degrees of freedom, and the critical value for a left-tailed test is "t_c = -1.76131."

The rejection region for this left-tailed test is "R = \{t: t <-1.76131\}"


3.The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{25-25}{10/\sqrt{15}}=0"



4. Since it is observed that "t =0>-1.76131= t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=14" degrees of freedom, "t=0" is "p = 0.5," and since "p= 0.5>0.05=\alpha," it is concluded that the null hypothesis is not rejected.


5. Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 25, at the "\alpha = 0.05" significance level.